prettydevil45:

Click an item in the list or group of pictures at the bottom of the problem and, holding the button down, drag it into the correct position in the answer box. Release your mouse button when the item is place. If you change your mind, drag the item to the trashcan. Click the trashcan to clear all your answers. Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (7, 2) (5, 4)

memeXpapi:

k

Katrin369:

didn't I answer this yesterday?

prettydevil45:

no

Katrin369:

what do you think is the answer?

prettydevil45:

1/7

prettydevil45:

y=mx+b

Katrin369:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @prettydevil45 y=mx+b \(\color{#0cbb34}{\text{End of Quote}}\) thats the form they gave you now use the form and put your answers there

prettydevil45:

ok y= !/7x+3

Katrin369:

how did you get that?

prettydevil45:

idk just guessed

Katrin369:

why is there a question mark?

prettydevil45:

idk

Katrin369:

ok i will explained it to youbut with different letters so you can do it and hopefully understand.

prettydevil45:

k

Katrin369:

1) The standard form of the equation of the line passing through two points is: Ax + By = C 2) You can find the equation of al line given two points in a two stages process: i) Find the slope: slope = m = [y₂ - y₁] / [ x₂ - x₁] = [1 - 2] / [5 - (-2) ] = - 1 / 7

Katrin369:

ii) set the point slope equation: y - y1 = m [x - x1] y - 1 = (-1/7) (x - 5) -> 7 ( y - 1) = - (x - 5) -> 7y - 7 = - x + 5 -> x + 7y = 5 + 7 -> x + 7y = 12

Katrin369:

hope this helps

prettydevil45:

sorry it dones not

prettydevil45:

@kaleighh533 can u help me

prettydevil45:

@Mathmachine2021

kaleighh533:

@prettydevil45 i dunno how to do this one

prettydevil45:

mann

kaleighh533:

@SkilerLuke or @MiraAngel

SkilerLuke:

I don't know im sorry

MiraAngel:

y=mx+b m: slope b: y-intercept (7, 2) (5, 4) Slope: (y2-y1)/(x2-x1) (x,y) (4-2)/(5-7) 2/-2 -1 y-intercept: (see graph) When x is 0, y is 9. 9 y=mx+b y=-1x+9

1 attachment